OpenStudy (anonymous):

Please help: 4/5x + 5 < -3

OpenStudy (anonymous):

I haven't done many problems like this before, what part should I do first?

OpenStudy (anonymous):

start by subtracting 5 from both sides so that you get the x by itself

OpenStudy (anonymous):

Ok so 4/5x < 2

OpenStudy (anonymous):

actually, doing it that way is going to make it more complicated than it needs to be ... one moment

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

actually, that ended up being the quickest way ... but subtracting from both sides leaves you with 4/5x < -8

OpenStudy (anonymous):

because -3 -5 = -8

OpenStudy (anonymous):

;P I always get mixed up with that even though I've gone over it a million times

OpenStudy (anonymous):

ok, right now with the equation like it is (4/5x < -8) ... you know "one" of the answers is zero because, well just because ... I forget why I just know it's true so just trust me and someone else who's taken math in the past decade will chime in in a bit to tell you

OpenStudy (anonymous):

ok so what should you do next?

OpenStudy (anonymous):

But, because they're tricky there's another answer too!

OpenStudy (anonymous):

next, multiply both sides by 5x to get that 5x out of the denominator

OpenStudy (anonymous):

ok so it would be 4x < -40 right

OpenStudy (anonymous):

4 = -40x

OpenStudy (jiteshmeghwal9):

\[4 < -3(5x+5)\]

OpenStudy (anonymous):

because \[\frac{ 4(5x) }{ 5x } = 4\] and \[-8 (5x )= -40x\]

OpenStudy (anonymous):

It's important to note, if this is (4/5)x + 5 < -3 or 4/(5x) + 5 < -3 These are not the same.

OpenStudy (jiteshmeghwal9):

\[4<-15x-15\]

OpenStudy (jiteshmeghwal9):

4+15<-15x

OpenStudy (anonymous):

@californiagirl55 my last equation should have had 4 < -40x (with < not =)

OpenStudy (jiteshmeghwal9):

19<-15x

OpenStudy (jiteshmeghwal9):

19/-15<x

OpenStudy (anonymous):

@qpHalcy0n has said the smartest thing in here thusfar ... I've been solving it as if the equation was \[\frac{ 4 }{ 5x } + 5 < 3\] And, @jiteshmeghwal9 is obviously answering it as if the question was \[\frac{ 4 }{5x + 5 } < 3\] which are completely different things

OpenStudy (anonymous):

grrr except I left off the negatives on the threes

OpenStudy (jiteshmeghwal9):

so what is the right question @RoamingBlue ??????/

OpenStudy (anonymous):

I have no idea *grins* wasn't my question to start with :D

OpenStudy (anonymous):

But if you solve it going down the path I was on you end up with 4 < -40x then 4/-40 < x or -1/10 < x < 0 as the final solution